OpenStudy (anonymous):

Which of the following is a linear function? (1 point) x = y – 3 y = x3 + 2 y = 6x2 y =

OpenStudy (anonymous):

last one seems to be missing something

OpenStudy (anonymous):

d: y=2/3x

OpenStudy (anonymous):

A) \[y=\frac{2}{3x}\] or B) \[y=\frac{2}{3}x\]

OpenStudy (anonymous):

the first one you can write at \(y=x+3\) which is actually not a function, it is an equation, but no matter this question was written by an idiot can't help that

OpenStudy (anonymous):

so it is in fact a line not sure what D though

OpenStudy (anonymous):

D: Y= 2/3x

OpenStudy (anonymous):

i cant figure it out

OpenStudy (anonymous):

go with A

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

by question about D is this: is the x in the denominator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then definitely go with A

OpenStudy (anonymous):

ok thanks!